# SAFREE: Training-Free and Adaptive Guard for Safe Text-to-Image And Video Generation

### v1

@@ 1 Introduction
Recent advancements in Generative AI have significantly impacted various modalities, including text (Brown, 2020), code (Chen et al., 2021), audio (Kreuk et al., 2022; Copet et al., 2024), image (Podell et al., 2023; Ho et al., 2022), and video generation (Ho et al., 2022; Kondratyuk et al., 2023; Yoon et al., 2024; openai, 2024). Generation tools such as DALL·E 3, Midjourney, Sora, and KLING have seen significant growth, enabling a wide range of applications in digital art, AR/VR, and educational content creation. However, these tools/models also carry the risk of generating content with unsafe concepts such as bias, discrimination, sex, or violence. Moreover, the definition of “unsafe content” varies according to societal perceptions. For example, individuals with Post-Traumatic Stress Disorder (PTSD) might find specific images (e.g., skyscrapers, deep-sea scenes) distressing. This underscores the need for an adaptable, flexible solution to enhance the safety of generative AI while considering individual sensitivities.
@@
This paper presents SAFREE, a training-free, adaptive plug-and-play mechanism for any diffusion-based generative model to ensure safe generation without altering well-trained model weights. SAFREE employs unsafe concept filtering in both textual prompt embedding and visual latent space, thereby enhancing the fidelity, quality, and efficiency for safe visual content generation. Specifically, SAFREE first identifies the unsafe concept subspace, i.e., the subspace within the input text embedding space that corresponds to undesirable concepts, by concatenating the column vectors of unsafe keywords. Then, to measure the proximity of each input prompt token to the unsafe/toxic subspace, we mask each token in the prompt and calculate the projected distance of the masked prompt embedding to the subspace. Given the proximity, SAFREE aims to filter out tokens that drive the prompt embedding closer to the unsafe subspace. Rather than directly removing or replacing unsafe tokens—which can compromise the coherence of the input prompt and degrade generation quality—SAFREE efficiently projects the identified unsafe tokens into a space orthogonal to the unsafe concept subspace while maintaining them on the original input embedding space. Such orthogonal projection design aims to preserve the overall integrity and the safe content within the original prompt while filtering out harmful content in the embedding space. To balance the trade-off between filtering out toxicity and preserving safe concepts (examples shown in Fig. 1), SAFREE incorporates a novel self-validating filtering scheme, which dynamically adjusts the number of denoising steps for applying filtered embeddings, enhancing the suppression of undesirable prompts when needed. Additionally, since we observe that unsafe content usually emerges at the regional pixel level, SAFREE extends filtering to the pixel space using a novel adaptive latent re-attention mechanism within the diffusion latent space. It selectively reduces the influence of features tied to the detected unsafe prompt, ensuring desirable outputs without drawing attention to those content. In the end, SAFREE filters out unsafe content simultaneously in both the embedding and pixel spaces. This approach ensures flexible and adaptive safe T2I/T2V generation that efficiently handles a broad range of unsafe concepts without extra training or modifications to model weights, while preserving the quality of safe outputs.
@@
Empirically, SAFREE achieves the state-of-the-art performance on five popular safe T2I benchmarks (I2P (Schramowski et al., 2023), P4D (Chin et al., 2024), Ring-A-Bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024b), and UnlearnDiff (Zhang et al., 2023b)) compared to other training-free methods, while also demonstrating competitive results against top training-based methods along with better efficiency, lower resource requirements and better inference-time adaptability across diverse unsafe concepts. We further apply our SAFREE to various T2I diffusion backbones (e.g., SDXL (Podell et al., 2023), SD-v3 (stabilityai, 2024)) and T2V models (ZeroScopeT2V (zeroscope, 2024), CogVideoX (Yang et al., 2024d)), showcasing SAFREE’s strong generalization and flexibility by effectively managing unsafe concept outputs across different models and tasks. As generative AI evolves rapidly, our training-free safeguard, SAFREE, sets a new strong baseline for generative AI safety, ensuring safe and ethical generative practices across diverse applications, from image to video synthesis, meeting the growing needs of the AI community.
@@
  • We propose SAFREE, a strong, adaptive, and training-free safeguard for text-to-image and text-to-video generation that ensures safety by jointly filtering out unsafe concepts in both textual embedding and visual latent spaces through conceptual proximity analysis.
  • SAFREE achieves state-of-the-art performance among training-free methods for concept removal in visual generation while maintaining high-quality outputs for desirable concepts, and it exhibits competitive results compared to training-based methods.
@@ 2.1 T2I attacks
Recent works address vulnerabilities in generative models, including LLMs (Patil et al., 2023; Liu et al., 2024b), VLMs (Zhao et al., 2023), and T2I models (Yang et al., 2024c; Wang et al., 2024; Li et al., 2024c). Cross-modality jailbreaks, like Shayegani et al. (2023), pair adversarial images with prompts to disrupt VLMs without accessing the language model. Tools like Ring-A-Bell (Tsai et al., 2024) and automated frameworks by Kim et al. (2024b) and Li et al. (2024a) focus on model-agnostic red-teaming and adversarial prompt generation, revealing safety flaws. Methods by Ma et al. (2024), Yang et al. (2024b), and Mehrabi et al. (2023) exploit text embeddings and multimodal inputs to bypass safeguards, using strategies like adversarial prompt optimization and in-context learning (Chin et al., 2024; Liu et al., 2024c). These works highlight vulnerabilities in T2I models.
@@ 2.2 Safe T2I generation
Training-based: Training-based approaches (Lyu et al., 2024; Pham et al., 2024; Zhang et al., 2024) ensure safe T2I generation by removing unsafe elements, as in Li et al. (2024c) and Gandikota et al. (2023), or using negative guidance. Adversarial training frameworks like Kim et al. (2024a) neutralize harmful text embeddings, while works like Das et al. (2024) and Park et al. (2024) filter harmful representations through concept removal and preference optimization. Fine-tuning methods such as Lu et al. (2024)’s cross-attention refinement and Heng & Soh (2023)’s continual learning remove inappropriate content. Latent space manipulation, explored by Liu et al. (2024a) and Li et al. (2024b), enhances safety using self-supervised learning. While effective, they require extensive fine-tuning, degrade image quality, and lack inference-time adaptation. Our SAFREE is training-free, dynamically adapts to concepts, and controls filtering strength w/o modifying weights, offering efficient safety across T2I and T2V models.
Training-free: Training-free methods for safe T2I generation adjust model behavior without retraining. These include (1) Closed-form weight editing, like Gandikota et al. (2024)’s model projection editing and Gong et al. (2024)’s target embedding methods, which remove harmful content while preserving generative capacity, and Orgad et al. (2023)’s minimal parameter updates to diffusion models. (2) Non-weight editing, such as Schramowski et al. (2023)’s Safe Latent Diffusion using classifier-free guidance and Cai et al. (2024)’s prompt refinement framework. However, these methods lack robustness and test-time adaptation. Our method is training-free, dynamically adjusts filtering based on prompts, and extends to other architectures and video tasks without weight edits, offering improved scalability and efficiency.
@@ 3 SAFREE: Training-Free and Adaptive Guard for Safe Text-to-Image and Video Generation
Recent approaches (Gandikota et al., 2024; Gong et al., 2024; Lu et al., 2024; Chavhan et al., 2024; Yang et al., 2024a) have demonstrated the effectiveness of weight modification through unlearning or model editing to prevent the generation of harmful (e.g., pornography, self-harm, violence), biased (e.g., racial or social stereotypes, ageism), or otherwise undesirable (e.g., public, copyrighted) visual content in text-to-image generation models. However, these methods have limited flexibility because they 1) require storing individual model weights for each concept to be removed, 2) inherently reduce the backbone model’s generative capabilities through unlearning, and 3) necessitate distinct solutions for safe generation across different models (i.e., modified model weights). To address this, we propose SAFREE, a training-free, yet adaptive remedy for safe T2I and T2V generation (See Fig. 2). We first determine the trigger tokens that can potentially induce toxicity based on the proximity between masked input prompt embeddings and toxic concept subspace (Sec. 3.1). Next, we propose projecting the detected undesirable tokens’ embeddings into a space orthogonal to the toxic concept subspace, while ensuring they remain within the input embedding space (Sec. 3.2). Additionally, we introduce an adaptive re-attention strategy, called self-validating filtering, in the latent space during the de-noising process, facilitating a robust joint filtering mechanism across both text and visual embedding spaces (Sec. 3.3). Finally, we extend our approach to high-resolution models like SDXL (Podell et al., 2023), DiT-based image diffusion (SD-v3) (Peebles & Xie, 2023), and representative text-to-video generative models such as ZeroScopeT2V (zeroscope, 2024) and CogVideoX-5B (Yang et al., 2024d) (Sec. 3.4).
Figure 2: Illustration of SAFREE framework. Based on proximity analysis between the masked token embeddings and the toxic subspace $\mathcal{C}$, we detect unsafe tokens and project them into orthogonal to the toxic concept (in red), but still be in the input space $\mathcal{I}$ (in green). SAFREE adaptively controls the filtering strength in an input-dependent manner, which also regulates a latent-level re-attention mechanism. Note that our approach can be broadly applied to various image and video diffusion backbones.
@@ 3.1 Adaptive Token Selection based on Toxic Concept Subspace Proximity
where $\omega$ is a hyperparameter controlling the guidance scale. $\bm{p}$ and $\emptyset$ denote the embedding of the input prompt and null text, respectively. The negative prompt is applied by replacing $\emptyset$ with the embedding of the negative prompt. We note that even if the input prompts are adversarial and unreadable to humans, such as those generated by adversarial attack methods through prompt optimization (Tsai et al., 2024; Yang et al., 2024b; Chin et al., 2024), they are still encoded within the same text embedding space, like CLIP (Radford et al., 2021), highlighting the importance and necessity of the feature embedding level unsafe concepts identification for safeguard. To this end, we propose to detect token embeddings that trigger inappropriate image generation and transform them to be distant from the toxic concept subspace: $\mathcal{C}=\left[\bm{c}_{0};\bm{c}_{1};...;\bm{c}_{K-1}\right]\in\mathbb{R}^{D\times K}$, which represents the embedding matrix that denotes the toxic subspace, where each column vector $\bm{c}_{k}$ corresponds to the embedding of the relevant text associated with the user-defined toxic concept, such as Sexual Acts or Pornography for Nudity concept.
To assess the relevance of tokens in the input prompt to the toxic concept subspace, we design a pooled input embedding $\overline{\bm{p}}_{\backslash i}\in\mathbb{R}^{D}$ that averages the token embeddings in $\bm{p}$ while masking out the $i$-th token. Suppose $\bm{z}\in\mathbb{R}^{K}$ be the vector of coefficients (i.e., the projection coordinates) of $\overline{\bm{p}}_{\backslash i}$ onto $\mathcal{C}$, it satisfies,
@@
We estimate the conceptual proximity of a token in the input prompt with $\mathcal{C}$ by computing the distance between the pooled text embedding obtained after masking out (i.e., removing) the corresponding token and $\mathcal{C}$. The residual vector $\bm{d}_{\backslash i}$, which is a component of $\overline{\bm{p}}_{\backslash i}$, is orthogonal to the subspace $\mathcal{C}$ is then formulated as follows:
@@
and $\bm{I}\in\mathbb{R}^{D\times D}$ denotes the identity matrix (See Fig. 2 middle left).
Refer to caption Refer to caption
A longer residual vector distance indicates that the removed token in the prompt is more strongly associated with the concept we aim to eliminate. In the end, we derive a masked vector $\bm{m}\in\mathbb{R}^{N}$ (where $N$ denotes the token length) to identify tokens related to the target concept, allowing us to subtly project them within the input token subspace while keeping them distant from the toxic concept subspace. We obtain a set of distances of masked token embeddings $\overline{\bm{p}}_{\backslash i},i\in[0,N-1]$ to the concept subspace, $D(\bm{p}|\mathcal{C})$, and select tokens for masking by evaluating the disparity between each token’s distance and the average distance of the set, excluding the token itself:
@@ 3.2 Safe Generation via Concept Orthogonal Token Projection
where $\odot$ indicates an element-wise multiplication operator. While our approach so far adaptively controls the number of token embeddings to be updated, it sometimes lacks flexibility in preserving the original generation capabilities for content outside the target concept. Recent observations (Kim et al., 2024a; Ban et al., 2024a) suggest that different denoising timesteps in T2I models contribute unevenly to generating toxic or undesirable content. Based on this insight, we propose a self-validating filtering mechanism during the denoising steps of the diffusion model that automatically adjusts the number of denoising timesteps conditioned on the obtained embedding (middle in Fig. 2). This mechanism amplifies the model’s filtering capability when the input prompt is deemed undesirable, while approximating the original backbone model’s generation for safe content. In the end, our updated input text embedding $\bm{p}_{safree}$ at a different denoising step $t$ is determined as follows:
$\begin{split}t^{\prime}&=\gamma\cdot\text{sigmoid}(1-\text{cos}(\bm{p},\bm{p}_{proj})),\\ \bm{p}_{safree}&=\begin{cases}\bm{p}_{safe}&\text{if}~{}~{}t\leq\text{round}(t^{\prime}),\\ \bm{p}&\text{otherwise},\end{cases}\end{split}$ (6)

### v2

@@ 1 Introduction
Recent advancements in Generative AI have significantly impacted various modalities, including text (Brown, 2020; Dubey et al., 2024), code (Chen et al., 2021; Liu et al., 2024a; Zhong et al., 2024), audio (Kreuk et al., 2022; Copet et al., 2024), image (Podell et al., 2023; Tian et al., 2024), and video generation (Ho et al., 2022; Kondratyuk et al., 2023; Yoon et al., 2024; Yang et al., 2024c). Generation tools such as DALL·E 3 (OpenAI, 2023), Midjourney (Midjourney, 2024), Sora (openai, 2024), and KLING (Kuaishou, 2024) have seen significant growth, enabling a wide range of applications in digital art, AR/VR, and educational content creation. However, these tools/models also carry the risk of generating content with unsafe concepts such as bias, discrimination, sex, or violence. Moreover, the definition of “unsafe content” varies according to societal perceptions. For example, individuals with Post-Traumatic Stress Disorder (PTSD) might find specific images (e.g., skyscrapers, deep-sea scenes) distressing. This underscores the need for an adaptable, flexible solution to enhance the safety of generative AI while considering individual sensitivities.
@@
This paper presents SAFREE, a training-free, adaptive plug-and-play mechanism for any diffusion-based generative model to ensure safe generation without altering well-trained model weights. SAFREE employs unsafe concept filtering in both textual prompt embedding and visual latent space, thereby enhancing the fidelity, quality, and efficiency for safe visual content generation. Specifically, SAFREE first identifies the unsafe concept subspace, i.e., the subspace within the input text embedding space that corresponds to undesirable concepts, by concatenating the column vectors of unsafe keywords. Then, to measure the proximity of each input prompt token to the unsafe/toxic subspace, we mask each token in the prompt and calculate the projected distance of the masked prompt embedding to the subspace. Given the proximity, SAFREE aims to filter out tokens that drive the prompt embedding closer to the unsafe subspace. Rather than directly removing or replacing unsafe tokens—which can compromise the coherence of the input prompt and degrade generation quality—SAFREE efficiently projects the identified unsafe tokens into a space orthogonal to the unsafe concept subspace while maintaining them on the original input embedding space. Such orthogonal projection design aims to preserve the overall integrity and the safe content within the original prompt while filtering out harmful content in the embedding space. To balance the trade-off between filtering out toxicity and preserving safe concepts (examples shown in Fig. 1), SAFREE incorporates a novel self-validating filtering scheme, which dynamically adjusts the number of denoising steps for applying filtered embeddings, enhancing the suppression of undesirable prompts only when needed. Additionally, since we observe that unsafe content usually emerges at the regional pixel level, SAFREE extends filtering to the pixel space using a novel adaptive latent re-attention mechanism within the diffusion latent space. It selectively reduces the influence of features in the frequency domain tied to the detected unsafe prompt, ensuring desirable outputs without drawing attention to those contents. In the end, SAFREE filters out unsafe content simultaneously in both the embedding and diffusion latent spaces. This approach ensures flexible and adaptive safe T2I/T2V generation that efficiently handles a broad range of unsafe concepts without extra training or modifications to model weights, while preserving the quality of safe outputs.
@@
Empirically, SAFREE achieves the state-of-the-art performance on five popular T2I benchmarks (I2P (Schramowski et al., 2023), P4D (Chin et al., 2024), Ring-A-Bell (Tsai et al., 2024), MMA-Diffusion (Yang et al., 2024a), and UnlearnDiff Zhang et al. (2023b)) outperforming other training-free safeguard methods with superior efficiency, lower resource use, and better inference-time adaptability. We further apply SAFREE to various T2I diffusion backbones (e.g., SDXL (Podell et al., 2023), SD-v3 (stabilityai, 2024)) and T2V models (ZeroScopeT2V (zeroscope, 2024), CogVideoX (Yang et al., 2024c)), showcasing SAFREE’s strong generalization and flexibility by effectively managing unsafe concept outputs across different models and tasks. As generative AI advances, SAFREE establishes a strong baseline for safety, promoting ethical practices across applications like image and video synthesis to meet the AI community’s needs.
@@
  • We propose SAFREE, a strong, adaptive, and training-free safeguard for T2I and T2V generation. This ensures more reliable and responsible visual content creation by jointly filtering out unsafe concepts in textual embeddings and visual latent spaces with conceptual proximity analysis.
  • SAFREE achieves state-of-the-art performance among training-free methods for concept removal in visual generation while maintaining high-quality outputs for desirable concepts, and it exhibits competitive results compared to training-based methods while maintaining better visual quality.
@@ 2.1 T2I attacks
Recent works address vulnerabilities in generative models, including LLMs (Zou et al., 2023; Patil et al., 2023; Wei et al., 2023; Liu et al., 2024c), VLMs (Zhao et al., 2023; Zong et al., 2024; Patil et al., 2024), and T2I models (Yang et al., 2024b; Wang et al., 2024; Li et al., 2024c). Cross-modality jailbreaks, like Shayegani et al. (2023), pair adversarial images with prompts to disrupt VLMs without accessing the language model. Tools like Ring-A-Bell (Tsai et al., 2024) and automated frameworks by Kim et al. (2024b) and Li et al. (2024a) focus on model-agnostic red-teaming and adversarial prompt generation, revealing safety flaws. Methods by Ma et al. (2024), Yang et al. (2024a), and Mehrabi et al. (2023) exploit text embeddings and multimodal inputs to bypass safeguards, using strategies like adversarial prompt optimization and in-context learning (Chin et al., 2024; Liu et al., 2024d). These works highlight vulnerabilities in T2I models.
@@ 2.2 Safe T2I generation
Training-based: Training-based approaches (Lyu et al., 2024; Pham et al., 2024; Zhang et al., 2024) ensure safe T2I generation by removing unsafe elements, as in Li et al. (2024c) and Gandikota et al. (2023), or using negative guidance. Adversarial training frameworks like Kim et al. (2024a) neutralize harmful text embeddings, while works like Das et al. (2024) and Park et al. (2024) filter harmful representations through concept removal and preference optimization. Fine-tuning methods such as Lu et al. (2024)’s cross-attention refinement and Heng & Soh (2023)’s continual learning remove inappropriate content. Latent space manipulation, explored by Liu et al. (2024b) and Li et al. (2024b), enhances safety using self-supervised learning. While effective, they require extensive fine-tuning, degrade image quality, and lack inference-time adaptation. SAFREE is training-free, dynamically adapts to concepts, and controls filtering strength without modifying weights, offering efficient safety across T2I and T2V models.
Training-free: Training-free methods for safe T2I generation adjust model behavior without retraining. These include (1) Closed-form weight editing, like Gandikota et al. (2024)’s model projection editing and Gong et al. (2024)’s target embedding methods, which remove harmful content while preserving generative capacity, and Orgad et al. (2023)’s minimal parameter updates to diffusion models. (2) Non-weight editing, such as Schramowski et al. (2023)’s Safe Latent Diffusion using classifier-free guidance and Cai et al. (2024)’s prompt refinement framework. However, these methods lack robustness and test-time adaptation. Our training-free method dynamically adjusts filtering based on prompts, and extends to other architectures and video tasks without weight edits, offering improved scalability and efficiency.
@@ 3 SAFREE: Training-Free and Adaptive Guard for Safe Text-to-Image and Video Generation
We propose SAFREE, a training-free, yet adaptive remedy for safe T2I and T2V generation (Fig. 2). We first determine the trigger tokens that can potentially induce toxicity based on the proximity between masked input prompt embeddings and toxic concept subspace (Sec. 3.1). The detected trigger token embeddings are projected onto a subspace orthogonal to the toxic concept subspace, while maintaining them within the input space (Sec. 3.2). SAFREE automatically adjusts the number of denoising timesteps conditioned on the text inputs through a self-validating filtering mechanism (Sec. 3.3). Additionally, we introduce an adaptive re-attention strategy in the latent space during the de-noising process, facilitating a robust joint filtering mechanism across both text and visual embedding spaces (Sec. 3.4). Finally, we extend our approach to high-resolution models like SDXL (Podell et al., 2023), DiT-based image diffusion (SD-v3) (Peebles & Xie, 2023), and representative text-to-video generative models such as ZeroScopeT2v (zeroscope, 2024) and CogVideoX (Yang et al., 2024c) (Sec. 3.5).
Figure 2: SAFREE framework. Based on proximity analysis between the masked token embeddings and the toxic subspace $\mathcal{C}$, we detect unsafe tokens and project them into orthogonal to the toxic concept (in red), but still be in the input space $\mathcal{I}$ (in green). SAFREE adaptively controls the filtering strength in an input-dependent manner, which also regulates a latent-level re-attention mechanism. Note that our approach can be broadly applied to various image and video diffusion backbones.
@@ 3.1 Adaptive Token Selection based on Toxic Concept Subspace Proximity
where $\omega$ is a hyperparameter controlling the guidance scale. $\bm{p}$ and $\emptyset$ denote the embedding of the input prompt and null text, respectively. The negative prompt is applied by replacing $\emptyset$ with the embedding of the negative prompt. We note that even if the input prompts are adversarial and unreadable to humans, such as those generated by adversarial attack methods through prompt optimization (Tsai et al., 2024; Yang et al., 2024a; Chin et al., 2024), they are still encoded within the same text embedding space, like CLIP (Radford et al., 2021), highlighting the importance and necessity of the feature embedding level unsafe concepts identification for safeguard. To this end, we propose to detect token embeddings that trigger inappropriate image generation and transform them to be distant from the toxic concept subspace: $\mathcal{C}=\left[\bm{c}_{0};\bm{c}_{1};...;\bm{c}_{K-1}\right]\in\mathbb{R}^{D\times K}$, which represents the embedding matrix that denotes the toxic subspace, where each column vector $\bm{c}_{k}$ corresponds to the embedding of the relevant text associated with the $k$-th user-defined toxic concept, such as Sexual Acts or Pornography for Nudity concept.
Detecting Trigger Tokens Driving Toxic Outputs. To assess the relevance of specific tokens in the input prompt to the toxic concept subspace, we design a pooled input embedding $\overline{\bm{p}}_{\backslash i}\in\mathbb{R}^{D}$ that averages the token embeddings in $\bm{p}$ while masking out the $i$-th token. Suppose $\bm{z}$$\in\mathbb{R}^{K}$ be the vector of coefficients (i.e., the projection coordinates) of $\overline{\bm{p}}_{\backslash i}$ onto $\mathcal{C}$, it satisfies,
@@
We estimate the conceptual proximity of a token in the input prompt with $\mathcal{C}$ by computing the distance between the pooled text embedding obtained after masking out (i.e., removing) the corresponding token and $\mathcal{C}$. The residual vector $\bm{d}_{\backslash i}$, which is the component of $\overline{\bm{p}}_{\backslash i}$ orthogonal to the subspace $\mathcal{C}$ is then formulated as follows:
@@
and $\bm{I}\in\mathbb{R}^{D\times D}$ denotes the identity matrix (See Fig. 2 middle left). A longer residual vector distance indicates that the removed token in the prompt is more strongly associated with the concept we aim to eliminate. In the end, we derive a masked vector $\bm{m}\in\mathbb{R}^{N}$ (where $N$ denotes the token length) to identify tokens related to the target concept, allowing us to subtly project them within the input token subspace while keeping them distant from the toxic concept subspace. We obtain a set of distances of masked token embeddings $\overline{\bm{p}}_{\backslash i},i\in[0,N-1]$ to the concept subspace, $D(\bm{p}|\mathcal{C})$, and select tokens for masking by evaluating the disparity between each token’s distance and the average distance of the set, excluding the token itself:
@@ 3.2 Safe Generation via Concept Orthogonal Token Projection
where $\odot$ indicates an element-wise multiplication operator. That is, for the $i$-th token, we use the projected safe embeddings $\bm{p}_{proj,i}$ only if it is detected as a toxic token ($m_{i}\odot\bm{p}_{proj,i},m_{i}=1$); otherwise, we retain the original (safe) token embeddings $\bm{p}_{i}$, since $(\bm{1}-\bm{m_{i}})\odot\bm{p}_{i},m_{i}=0$.
3.3 Adaptive Control of Safeguard Strengths with Self-validating Filtering
While our approach so far adaptively controls the number of token embeddings to be updated, it sometimes lacks flexibility in preserving the original generation capabilities for content outside the target concept. Recent observations (Kim et al., 2024a; Ban et al., 2024a) suggest that different denoising timesteps in T2I models contribute unevenly to generating toxic or undesirable content. Based on this insight, we propose a self-validating filtering mechanism during the denoising steps of the diffusion model that automatically adjusts the number of denoising timesteps conditioned on the obtained embedding (middle in Fig. 2). This mechanism amplifies the model’s filtering capability when the input prompt is deemed undesirable, while approximating the original backbone model’s generation for safe content. In the end, our updated input text embedding $\bm{p}_{safree}$ at a different denoising step $t$ is determined as follows:
$\begin{split}t^{\prime}=\gamma\cdot\text{sigmoid}(1-\text{cos}(\bm{p},\bm{p}_{proj})),\quad\quad\bm{p}_{safree}=\begin{cases}\bm{p}_{safe}&\text{if}~{}~{}t\leq\text{round}(t^{\prime}),\\ \bm{p}&\text{otherwise},\end{cases}\end{split}$ (6)